If the cost, C(x), for manufacturing x units of a certain product is given by C(x)= x^2 - 30x + 7000 find the number of units manufactured at a cost of $8000
Answer by richwmiller(17219)

x^2 - 30x + 7000=8000
x^2 - 30x -1000=0
(x-50) (x+20) = 0
x=50
x=-20
reject negative answers